OpenStudy (anonymous):

Write the double angle formula for cosine in two alternate forms by filling in the blanks using a trig identity or simplifying your answer at each step. a) cos(2t) = 1 - sin^2t = 1 ( "fill in the blank") = "fill in the blank" b) cos(2t) = 2cost^2(t) - 1 = 2cos^2(t) - ( "fill in the blank") = fill in the blank.

OpenStudy (anonymous):

Im trying to use the pythagorean identities but its always saying im wrong O.o

OpenStudy (anonymous):

@lgbasallote can you help :/

OpenStudy (lgbasallote):

this doesn't make sense... cos(2t) isn't equal to 1- sin^2 t...can you explain your question?

OpenStudy (anonymous):

2cos^2x-1

OpenStudy (anonymous):

Its an interactive website, ill take a screen shot wait

OpenStudy (anonymous):

Its some pre cal stuff just as a review. but even then i just cant seem to get it "-_-

OpenStudy (anonymous):

cos2t= 2cos^2 (x) -1 = 1- 2sin^2 (x)=cos^2 (x)-sin^2(x)

OpenStudy (lgbasallote):

that's a totally different thing

OpenStudy (anonymous):

I didnt write it down correctly? sigh im sorry

OpenStudy (anonymous):

oh i forgot the 2!!!

OpenStudy (anonymous):

the answer is already there,

OpenStudy (anonymous):

well what do you mean?

OpenStudy (anonymous):

the fill in the blank?

OpenStudy (anonymous):

i have already given the asnswer, just look at it to fill in the missing blanks